a shortbread recipe uses flour, butter, sugar and nuts in the ratio, by weight, of 9:6:3:2. How much of each ingredient is used in making 1kg of the mixture
Answer by pakhi(24)

The ratio is 9:6:3:2
If we add 9+6+3+2, we get the result as 20. So if 1kg of the mixture is divided into 20 parts, 9 parts of it will be flour,6 parts will be butter, 3 parts will be sugar and 2 parts will be nuts.
1kg=1000grams
Therefore (9/20)*1000=450g is flour
(6/20)*1000=300g is butter
(3/20)*1000=150g is sugar
(2/20)*1000=100g is nuts
check: 450+300+150+100=1000